OpenStudy (anonymous):

Solve the system by using a matrix equation. 2x-3y=2 7x-5y=7

OpenStudy (anonymous):

If we want: First find x should killing the y s ! Now starting together !:)

OpenStudy (anonymous):

What is your sejestion ?! I think we should have y s *5 .

OpenStudy (anonymous):

Now start : -5 .( 2x-3y=2) ___>-10x+15y =- 10 3. (7x-5y=7)_____>21x-15y=21 So we +ing allllll ! 11x=11 x=1 Understand ?!:) :D